If a sample of tap water contains 400 ppm of Ca2+ and 80 ppm of Mg2+, what is the hardness of this water sample in terms of equivalent concentration CaCO3, in mg/L?
Answer #1

MCaCo3/MCa=100/40=2.5....... as because the molar mass of calcium carbonate is 100 and atomic weight of Ca is 40

MCaCo3/MMg=100/24.3=4.1.......Molar mass of Magnesium is 24.3

We know,

Hardness= 2.5[Ca2+]+4.1[Mg2+] =2.5*400+4.1*80=1328 ppm

Hence the answer is 1328 ppm in CaCO3 scale and it is of high hardness.